# Wyckoff Transformer: Generation of Symmetric Crystals

### v3

@@ 2.1 Tokenization
Our work is based on the inductive bias that for stable materials space group symmetry and Wyckoff sites almost completely define the structure – more than 98% of the materials in MP-20 (Xie et al., 2021) and MPTS-52 (Baird et al., 2024) datasets, which together contain almost all experimentally stable structures from the Materials Project (Jain et al., 2013), have unique Wyckoff representations. Therefore it is safe to assume that for almost any Wyckoff representation there is either none, or just one stable material conforming to it. Symmetry captured by this discrete part is sufficient to determine properties of a material, such as piezoelectricity via non-centrosymmetry; direct/indirect band gap via positions of the valence/conduction bands in the Brillouin Zone, while the fractional coordinates can be linked to the magnitude of that property. We additionally prove this assumption by various predicting material properties, see section 3.2. Given a Wyckoff representation, coordinates can be determined as discussed in section 2.4.
We represent each structure as a set of tokens, as shown in Figure4. The first token contains the space group, the rest are divided into groups of three tokens, each representing a specific WP. The first token in each group is responsible for the type of atom that occupies the position following the site symmetry, while the last token is for the so-called enumeration. Several WPs can have the same site symmetry. To differentiate those WPs we enumerate them separately within each space group and site symmetry according to the conventional WP order (Aroyo et al., 2006). For example, in space group 225 present in Figure4 WP 4a is encoded as (m-3m, 0), 4b as (m-3m, 1), and 8c as (-43m, 0). A more comprehensive example can be found in Appendix S. The purpose of this encoding is to take advantage of the fact that, unlike Wyckoff letters, site symmetry definition is universal across different space groups. An ablation study comparing our representation with Wyckoff letters is in Appendix N.
@@ 2.1.1 Spherical harmonics
Enumerations are defined by an arbitrary convention, in this respect they are no better than Wyckoff letters. We address this with a representation that is defined consistently across space groups. Consider a Wyckoff position consisting of a set of $k$ symmetry operations $\{R_{i}\bm{x}+\bm{t}_{i},i=1...k\}$. We apply those operations to points $\bm{x}_{1}=[0,0,0]$ and $\bm{x}_{2}=[1,1,1]$ obtaining two matrices $W^{(1)}$ and $W^{(2)}$: $W_{i}^{(j)}=R_{i}\bm{x}_{j}+\bm{t}_{i}\bm{x}_{j}$. Finally, we convolve the transformed coordinates with spherical harmonics:
@@ 2.2 Model architecture
Elements, site symmetries, and enumerations are each embedded with a simple lookup table with trainable weights, the embeddings are concatenated; then we apply a linear layer. The reason is that we want to provide each head of the multihead attention with information from all three parts of a token.
Since our model is conditioned on space group, preventing data fragmentation is of utmost importance. To this end, the space group is not encoded just as a categorical variable. Building upon (AI4Science et al., 2023) and similarly to Levy et al. (2024) we use pyXtal to get one-hot-encoded $15\times 10$ matrix that represents symmetry elements on each axis for each space group, flatten it, discard the positions that do not vary across the dataset and use the resulting vector as the space group embedding. Then we apply a linear layer, so the representation becomes learnable – but still transferable between space groups.
@@
Property prediction We take the average of the Transformer encoder outputs tokens, excluding the token corresponding to the space group, compute a weighted average with weights being equal to the multiplicities of WPs, and use the result as input for a fully-connected neural network that outputs a scalar predicted value.
@@ 3.1.1 Datasets
We use MP-20 (Xie et al., 2021), which contains almost all experimentally stable materials in Materials Project (Jain et al., 2013) with a maximum of 20 atoms per unit cell, within 0.08 eV/atom of the convex hull, and formation energy smaller than 2 eV/atom, 45229 structures in total, split 60/20/20 into train, validation and test parts. Additionally, we train and evaluate WyFormer on MPTS-52 (Baird et al., 2024), a more challenging subset of Materials Projects containing materials with up to 52 atoms per unit cell.
@@ 3.1.2 Metrics
Novelty and uniqueness The purpose of de novo generation is to obtain new materials. Generating materials that already exist in the training dataset increases the model performance according to structure stability and similarity metrics, but such structures are useless for material design and just increase the gap between the proxy metrics and the model fitness for its purpose. Therefore we exclude generated materials that are not novel and unique from metric computation. On a deeper level, generative models for materials are subject to exploration/exploitation trade-off: the more physically similar are the sampled materials to the training dataset, the more likely they are stable and distributed similar to the data, but the less useful they are for the purpose of material design. From a purely machine learning point of view, novelty percentage serves a proxy metric for overfitting.
Stability determines whether the material, in fact, exists under normal conditions. It is estimated by computing energy above convex hull, and comparing it to a threshold. Materials Project is the source of the reference structures for the hull; The details are in Appendix G.
@@
– P1 is the percentage of the structures that have symmetry group P1. In MP-20 the corresponding number is just 1.7%, and yet more than a third of the structures generated by some state-of-the-art models lack symmetry beyond lattice translation. We argue that presence of symmetry is good proxy value for structure feasibility that is difficult to capture in standard DFT computations, and would require finite-temperature calculations and/or improved methodologies.
@@ 3.1.3 Methodology
WyCryst (Zhu et al., 2024) only supports a limited number of unique elements per structure, therefore we trained it on a subsection of MP-20 containing only binary and ternary compounds, 35575 in total. An evaluation of WyFormer trained on the same dataset is present in Appendix K. As WyCryst also produces Wyckoff representations, and not structures, the same CrySPR+CHGNet procedure was used to obtain them.
@@ 3.2 Material property prediction
CHGNet 34 MPTrj MP–20
WyFormer 25 234 MP–20
@@ Acknowledgements
This research/project is supported by the Ministry of Education, Singapore, under its Research Centre of Excellence award to the Institute for Functional Intelligent Materials (I-FIM, project No. EDUNC-33-18-279-V12). This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-028) and from the MAT-GDT Program at A*STAR via the AME Programmatic Fund by the Agency for Science, Technology and Research under Grant No. M24N4b0034. The computational work for this article was performed on resources at the National Supercomputing Centre of Singapore (NSCC). Computational work involved in this research work is partially supported by NUS IT’s Research Computing group. The research used computational resources provided by Constructor Tech.

### v4

@@ 2.1 Tokenization
Our work is based on the inductive bias that for stable materials space group symmetry and Wyckoff sites almost completely define the structure – more than 98% of the materials in MP-20 (Xie et al., 2021) and MPTS-52 (Baird et al., 2024) datasets, which together contain almost all experimentally stable structures from the Materials Project (Jain et al., 2013), have unique Wyckoff representations. Therefore, it is safe to assume that for almost any Wyckoff representation there is either none, or just one stable material conforming to it. Symmetry captured by this discrete part is sufficient to determine properties of a material, such as piezoelectricity via non-centrosymmetry; direct/indirect band gap via positions of the valence/conduction bands in the Brillouin Zone, while the fractional coordinates can be linked to the magnitude of that property. We additionally prove this assumption by predicting various material properties; see Section 3.2. Given a Wyckoff representation, coordinates can be determined as discussed in Section 2.4.
We represent each structure as a set of tokens, as shown in Figure4. The first token contains the space group; the rest are divided into groups of three tokens, each representing a specific WP. The first token in each group is responsible for the type of atom that occupies the position following the site symmetry, while the last token is for the so-called enumeration. Several WPs can have the same site symmetry. To differentiate those WPs we enumerate them separately within each space group and site symmetry according to the conventional WP order (Aroyo et al., 2006). For example, in space group 225 present in Figure4 WP 4a is encoded as (m-3m, 0), 4b as (m-3m, 1), and 8c as (-43m, 0). A more comprehensive example can be found in Appendix S. The purpose of this encoding is to take advantage of the fact that, unlike Wyckoff letters, site symmetry definition is universal across different space groups. An ablation study comparing our representation with Wyckoff letters is in Appendix N.
@@ 2.1.1 Spherical harmonics
Enumerations are defined by an arbitrary convention, in this respect they are no better than Wyckoff letters. We address this with a representation that is defined consistently across space groups. Consider a Wyckoff position consisting of a set of $k$ symmetry operations $\{R_{i}\bm{x}+\bm{t}_{i},i=1...k\}$. We apply these operations to points $\bm{x}_{1}=[0,0,0]$ and $\bm{x}_{2}=[1,1,1]$ obtaining two matrices $W^{(1)}$ and $W^{(2)}$: $W_{i}^{(j)}=R_{i}\bm{x}_{j}+\bm{t}_{i}\bm{x}_{j}$. Finally, we convolve the transformed coordinates with spherical harmonics:
@@ 2.2 Model architecture
Elements, site symmetries, and enumerations are each embedded with a simple lookup table with trainable weights, the embeddings are concatenated. Then we apply a linear layer to provide each head of the multihead attention with information from all three parts of a token.
Since our model is conditioned on space group, preventing data fragmentation is of utmost importance. To this end, the space group is not encoded just as a categorical variable. Building upon (AI4Science et al., 2023) and similarly to Levy et al. (2024) we use pyXtal to get one-hot-encoded $15\times 10$ matrix that represents symmetry elements on each axis for each space group, flatten it, discard the positions that do not vary across the dataset and use the resulting vector as the space group embedding. Then we apply a linear layer, so the representation becomes learnable — but still transferable between space groups.
@@
Property prediction We take the Transformer encoder outputs tokens, excluding the token corresponding to the space group, compute a weighted average with weights being equal to the multiplicities of WPs, and use the result as input for a fully-connected neural network that outputs a scalar predicted value.
@@ 3.1.1 Datasets
We use MP-20 (Xie et al., 2021), which contains almost all experimentally stable materials in Materials Project (Jain et al., 2013) with a maximum of 20 atoms per unit cell, within 0.08 eV/atom of the convex hull, and formation energy smaller than 2 eV/atom, 45 2294522945\,22945 229 structures in total, split 60/20/20 into train, validation and test parts. Additionally, we train and evaluate WyFormer on MPTS-52 (Baird et al., 2024), a more challenging subset of Materials Projects containing materials with up to 52 atoms per unit cell.
@@ 3.1.2 Metrics
Novelty and uniqueness The purpose of de novo generation is to obtain new materials. Generated materials that already exist in the training dataset increase the model performance according to structure stability and similarity metrics, but such structures are useless for material design and just increase the gap between the proxy metrics and the model fitness for its purpose. Therefore we exclude generated materials that are not novel and unique from metric computation. On a deeper level, generative models for materials are subject to exploration/exploitation trade-off: the more physically similar are the sampled materials to the training dataset, the more likely they are stable and distributed similar to the data, but the less useful they are for the purpose of material design. From a purely machine learning point of view, novelty percentage serves a proxy metric for overfitting.
Stability determines whether the material, in fact, exists under normal conditions. It is estimated by computing energy above convex hull, and comparing it to a threshold. Materials Project is the source of the reference structures for the hull. The details are in Appendix G.
@@
– P1 is the percentage of the structures that have symmetry group P1. In MP-20 the corresponding number is just 1.7%. We argue that presence of symmetry is good proxy value for structure feasibility that is difficult to capture in standard DFT computations, and would require finite-temperature calculations and/or improved methodologies.
@@ 3.1.3 Methodology
WyCryst (Zhu et al., 2024) only supports a limited number of unique elements per structure, therefore we trained it on a subset of MP-20 containing only binary and ternary compounds, 35 5753557535\,57535 575 in total. An evaluation of WyFormer trained on the same dataset is present in Appendix K. As WyCryst also produces Wyckoff representations, and not structures, the same CrySPR+CHGNet procedure was used to obtain them.
@@ 3.2 Material property prediction
CHGNet 34 MPTrj MP-20
WyFormer 25 234 MP-20
@@ Acknowledgements
This research/project is supported by the Ministry of Education, Singapore, under its Research Centre of Excellence award to the Institute for Functional Intelligent Materials (I-FIM, project No. EDUNC-33-18-279-V12). This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-028) and from the MAT-GDT Program at A*STAR via the AME Programmatic Fund by the Agency for Science, Technology and Research under Grant No. M24N4b0034. The computational work for this article was performed on resources at the National Supercomputing Centre of Singapore (NSCC). Computational work involved in this research work is partially supported by NUS IT’s Research Computing group. The research used computational resources provided by Constructor Tech. This research was supported in part through computational resources of HPC facilities at HSE University.